OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (- 3, - 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 6x − { Blank }

OpenStudy (anonymous):

@David27 Do you know how to do this?

OpenStudy (david27):

no sorry but he might know @ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (dan815):

x intercept of -8 means the point (-8,0) is a solution x int of 2 means the point (2,0) is a solution to your equation

OpenStudy (dan815):

y = x2 + 6x − { Blank } y = x2 + 6x −B knowing that (-8,0) is a point of this parabola so, plugging that in for x,y 0=(-8)^2+ 6*(-8) -B B=? solve for B you can also use the other point (2,0) 0=2^2+6*2-B B=?

OpenStudy (anonymous):

So solve 2^2+6*2 first?